### v1

# The LHC has ruled out Supersymmetry – really?

Abstract
Despite early hopes that the Large Hadron Collider (LHC) would quickly unveil supersymmetric particles, none have been detected to date. This review examines the impact of the LHC results on the viability of weak-scale supersymmetry, and discusses whether the possibility of discovering supersymmetric particles remains within reach.
keywords:
new physics , supersymmetry , LHC
PACS:
12.60.-i , 12.60.Jv , 14.80.Ly
journal: Nuclear Physics B
[…]
[a]organization=Laboratoire de Physique Subatomique et de Cosmologie (LPSC), Université Grenoble-Alpes, CNRS/IN2P3, postcode=F-38026, city=Grenoble, country=France
[…]
[b]organization=Université Claude Bernard Lyon 1, CNRS/IN2P3, Institut de Physique des 2 Infinis de Lyon, UMR 5822, postcode=F-69622, city=Villeurbanne, country=France
[…]
[c]organization=Theoretical Physics Department, CERN, addressline=CH-1211 Geneva 23, country=Switzerland
[…]
[d]organization=Institut Universitaire de France (IUF), postcode=F-75005, city=Paris, country=France
1 Introduction
Supersymmetry (SUSY) is a theoretical framework that extends the Standard Model (SM) by a new symmetry between bosons and fermions. In this scenario, every known particle (more precisely, every SM degree of freedom) has a supersymmetric partner, differing in spin by half a unit. SUSY has been the prime paradigm for new physics beyond the SM for decades. For a detailed introduction, see e.g. Martin (1998).
One of the key motivations is the so-called hierarchy problem: within the SM, quantum corrections drive the Higgs boson mass to extremely high values, requiring a fine-tuning to maintain it at the electroweak scale. Weak-scale SUSY naturally solves this issue by ensuring that the divergent corrections from fermionic and bosonic partners cancel out, thereby stabilising the Higgs mass. Additionally, SUSY predicts the unification of gauge couplings at high energy scales. Another compelling aspect of supersymmetric models is the ability to provide viable dark matter candidates, particularly in models where R-parity is conserved: in this case, the lightest supersymmetric particle (LSP), often assumed to be a neutralino, remains stable and could account for the observed dark matter density in the universe.
If R-parity is conserved, SUSY particle production at colliders is characterised by a distinctive experimental signature: SM particles accompanied by significant missing transverse energy (MET). The LSP, which remains stable under these conditions, naturally fulfills the role of a weakly interacting massive particle (WIMP) and can account for the observed dark matter abundance over a broad parameter space.
Early expectations for the discovery of SUSY at the LHC were high, particularly for strongly interacting superpartners. Theoretical arguments, but also global fits of constrained SUSY scenarios —such as the CMSSM— incorporating constraints from flavour physics, low-energy precision measurements, and dark matter relic density (see e.g. Allanach et al. (2007); Buchmueller et al. (2008); Trotta et al. (2008)), suggested that superparticle masses should be within reach of the LHC initial runs, typically below or around 1 TeV. However, searches by the ATLAS and CMS experiments (for a review, see e.g. Canepa (2019)) have ruled out substantial regions of this parameter space, challenging initial expectations. [^1] Comprehensive overviews of search results, including summary plots, are available at the ATLAS and CMS wiki pages https://twiki.cern.ch/twiki/bin/view/AtlasPublic/SearchesPublicResults and https://cms-results.web.cern.ch/cms-results/public-results/publications/SUS/index.html
Indeed, mass limits are being pushed higher and higher by the LHC experiments. The current bounds reported by ATLAS and CMS reach up to about $2.4$ TeV for gluinos and squarks of the first two generations, $1.3$ TeV for stops and sbottoms, and $1$ TeV for electroweak-inos [Kwon (2025)]. It has to be kept in mind, however, that these bounds are derived in the context of simplified models, assuming the presence of only few (typically 2–3) new particles and simple decay patterns (often a single decay channel). Realistic SUSY models can be much more complicated, with a large variety of relevant states and intricate mixing patterns affecting cross sections, decay branching ratios, and in specific cases even the lifetimes of the new states. Thus, even in the simplest realistic model, the so-called Mininal Supersymmetric Standard Model (MSSM), mass bounds can be weaker than those listed above.
Be as it may, the absence of conclusive evidence for supersymmetry at the LHC has led to growing frustration within the high-energy physics community. Many researchers nowadays question whether SUSY remains a viable solution to the shortcomings of the SM, and, in popular discourse, there is a growing perception that the LHC has effectively “ruled out” SUSY.
In this article, we address the question whether this claim is true or an oversimplification.
2 Higgs mass, hierarchy and naturalness problems
In order to break the electroweak symmetry and give masses to the $W$ and $Z$ bosons and the matter fermions, some scalar field must acquire a non-zero vacuum expectation value (VEV). In the SM, this field is elementary, leading to an elementary scalar “Higgs” boson of mass $m_{H}$. However, elementary scalar masses are notoriously unstable in the SM. For example, the one-loop diagram shown in Fig. 1(left) is quadratically divergent, yielding
$\delta m_{H}^{2}={\cal O}\left({\alpha\over 4\pi}\right)\,\Lambda^{2}\,,$ (1)
where the cutoff $\Lambda$ represents the scale up to which the SM remains valid; this may be $M_{X}\geq 10^{15}$ GeV in a grand unified theory (GUT), or the Planck scale $M_{\rm P}\sim 10^{19}$ GeV, if no new physics sets in before. Moreover, in GUTs, the couplings of a light Higgs boson to the heavy Higgs fields present in the theory also yield $\delta m_{H}^{2}={\cal O}(M_{X}^{2})$. This is in principle not a problem in a renormalisable theory. However, it requires an extreme fine-tuning of counterterms order by order in pertubation theory, and this seems unnatural.
Figure 1: One-loop corrections to $m_{H}^{2}$, due to a Dirac fermion $f$ (left), and a scalar $S$ (right); from Martin (1998).
These large corrections to the SM Higgs boson mass, which should be $m_{H}={\cal O}(m_{W})$, raise problems at two levels:
(i) to arrange, in the first place, for $m_{H}$ to be many orders smaller than other fundamental mass scales such as $M_{X}$ or $M_{\rm P}$ —the hierarchy problem, and (ii) to avoid corrections $\delta m_{H}^{2}$ to the Higgs boson mass squared, which are much larger than $m_{H}^{2}$ itself —the naturalness problem.
In SUSY, SM particles and their superpartners have the same couplings but differ in spin by half a unit. Since fermion (Fig. 1 left) and boson (Fig. 1 right) loop diagrams have opposite signs, the quadratic sensitivity to the cutoff $\Lambda$ in eq. (1) cancels. The remaining correction is
$\delta m_{H}^{2}={\cal O}\left({\alpha\over 4\pi}\right)\;\left|m_{f}^{2}-m_{S}^{2}\,\right|\,,$ (2)
where $m_{f}$ and $m_{S}$ stand for the masses of the SM particles and their superpartners. This is no larger than the physical value, $\delta m_{H}^{2}\lesssim m_{H}^{2}$, provided
[…]
This is (or was) the main motivation to expect SUSY to be realised at the TeV energy scale! The argument is in fact most severe for stops ($\tilde{t}$), the scalar partners of the top quark, since the largest contribution to $\delta m_{H}^{2}$ in the SM comes from top-quark loops. Therefore the fact that LHC searches seem to push $m_{\tilde{t}}\gtrsim 1$ TeV is perceived as a severe problem.
On the other hand, in SUSY, the Higgs boson mass is not a free parameter like in the SM. In the MSSM, the Higgs sector consists of two complex doublets $H_{u}=(H^{+}_{u},H^{0}_{u})^{T}$, $H_{d}=(H^{0}_{d},H^{-}_{d})^{T}$ with the scalar potential for the (neutral) Higgs fields given by
${\cal V}=$ $\>(|\mu|^{2}+m_{H_{u}}^{2})\,|H^{0}_{u}|^{2}+(|\mu|^{2}+m_{H_{d}}^{2})\,|H^{0}_{d}|^{2}-(bH_{u}^{0}H_{d}^{0}+{\text{h}.c.})$
$+{\textstyle\frac{1}{8}}(g_{1}^{2}+g_{2}^{2})\,(|H_{u}^{0}|^{2}-|H_{d}^{0}|^{2})^{2}\,.$ (4)
Here $\mu$ is a SUSY-conserving Higgs mixing parameter from the superpotential (${\cal W}\in\mu H_{u}H_{d}$), while $m_{H_{u}}^{2}$, $m_{H_{d}}^{2}$ and $b$ are soft SUSY-breaking parameters; see Martin (1998) for details. The structure is that of a two-Higgs doublet model of type II and leads to five physical Higgs states, a neutral CP-odd, two neutral CP-even, and two charged Higgs bosons denoted by $A^{0}$, $h^{0}$, $H^{0}$, and $H^{\pm}$, respectively. An important point here is that the quartic term in eq. (4) is a function of the gauge couplings, in contrast to being a free parameter in the SM. As a consequence, after electroweak symmetry breaking, there is an upper bound on the mass of the lighter of the two CP-even Higgs bosons $h^{0}$:
$m_{h}\leq m_{Z}\,|\cos 2\beta|$ (5)
at tree level. Here, $\tan\beta=v_{u}/v_{d}$ is the ratio of the VEVs at the minimum of the potential, $v_{u}\equiv\langle H_{u}^{0}\rangle$ and $v_{d}\equiv\langle H_{d}^{0}\rangle$, and $v^{2}\equiv v_{u}^{2}+v_{d}^{2}=2\,m_{Z}^{2}/(g_{1}^{2}+g_{2}^{2})$.
Figure 2: Dependence of the light MSSM Higgs boson mass, $M_{h}$, on the ratio $X_{t}/M_{S}$; from Arbey et al. (2012a).
Fortunately, $m_{h}$ is subject to large radiative corrections. Again, due to the large top Yukawa coupling, the largest effect comes from the top-stop sector. In the decoupling limit $m_{A}\gg m_{Z}$, one gets at the one-loop level
$m_{h}^{2}=m_{Z}^{2}\cos^{2}2\beta+\frac{3}{4\pi^{2}}\frac{m_{t}^{4}}{v^{2}}\left(\log\frac{M_{S}^{2}}{m_{t}^{2}}+\frac{X_{t}^{2}}{M_{S}^{2}}\left(1-\frac{X_{t}^{2}}{12\,M_{S}^{2}}\right)\right)\,.$ (6)
Here $M_{S}^{2}=m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}$ with $m_{\tilde{t}_{1,2}}$ the stop masses, and $X_{t}=A_{t}-\mu\cot\beta$ is the stop mixing parameter. To achieve the observed value of $m_{h}=125$ GeV [Aad et al. (2012); Chatrchyan et al. (2012)] either the average stop mass must be large, $M_{S}\equiv\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}\gtrsim 3$ TeV, or the stop mixing parameter $|X_{t}|$ must be around twice $M_{S}$, as illustrated in Fig. 2 (taken from Arbey et al. (2012a), who, together with Brummer et al. (2012), also discuss the implications for GUT models). [^2] There is a vast literature on this subject, and on the topic of precision calculations of $m_{h}$. For reviews, we refer the reader to [Djouadi (2008); Draper and Rzehak (2016)]. This leads to a tension with the expectation of light stops (and not too heavy gluinos, due to their loop contributions to stop masses) motivated by the hierarchy problem, and with the wish for small radiative corrections to satisfy naturalness arguments.
Models beyond the MSSM can alleviate the problem through extra tree-level contributions to $m_{h}^{2}$. This is the case, for instance, in models with extra singlet scalars, see e.g. [Ellwanger et al. (2010)], or when gauginos are Dirac instead of Majorana states [Benakli et al. (2013)]. Either way a tension remains and has led many people to doubt SUSY as a solution to the hierarchy and naturalness problems.
One way out is to abandon the hierarchy problem altogether and use gauge-coupling unification and dark matter as the only guiding principles, as proposed in Arkani-Hamed and Dimopoulos (2005); Giudice and Romanino (2004); Arkani-Hamed et al. (2005). In this case, all the MSSM scalars become (ultra)heavy, except for a single finely tuned SM-like Higgs boson. Yet, the fermionic superpartners (gauginos and higgsinos) can remain light, protected by chiral symmetry, and account for the successful unification of gauge couplings. This is known as Split SUSY and has distinctive consequences for LHC signatures.
An even more extreme approach is to consider that gauginos and higgsinos are also very heavy, with masses near the SUSY-breaking scale $M_{S}$. This defines the High-Scale SUSY framework Hall and Nomura (2010), where supersymmetry is effectively decoupled from the low-energy spectrum. In this scenario, one forgoes not only the naturalness argument but also the supersymmetric explanation for dark matter—since the LSP is no longer light—and the successful gauge coupling unification typically offered by SUSY. Nevertheless, supersymmetry leaves an imprint at low energies through the matching conditions between the high-scale SUSY theory and the effective Standard Model. In particular, the Higgs quartic coupling $\lambda$ retains information about the underlying SUSY structure. As a result, even when SUSY is broken at very high energies, it still predicts a relatively light Higgs boson, with its mass carrying indirect sensitivity to parameters like $M_{S}$ and $\tan\beta$.
The relation between the observed Higgs mass at 125 GeV and the SUSY mass scale in the Split SUSY and High-Scale SUSY scenarios was investigated by Giudice and Strumia (2012); Arbey et al. (2012a); Bagnaschi et al. (2014) and is illustrated in Fig. 3.333This may be strongly influenced by tree-level contributions to $\lambda$ from a non-minimal field content.
Figure 3: Predictions for the Higgs mass $m_{h}$ in Split SUSY (red) and in High-Scale SUSY (blue) for $\tan\beta=\{1,2,4,50\}$; from Giudice and Strumia (2012).
3 Little Hierarchy problem and electroweak naturalness
One can also adopt the more optimistic point of view of Baer et al., that SUSY offers a ’t Hooft technically natural solution to the (Big) Hierarchy problem, and what matters is the electroweak finetuning. This is related to the Little Hierarchy problem: why is there a gap between the weak scale $m_{weak}\sim m_{W,Z,h}\sim 100$ GeV and the apparently lowest-possible soft SUSY breaking scale $M_{S}>1-2$ TeV?
In the MSSM, the weak scale is related to the soft SUSY breaking terms and the $\mu$ parameter by the electroweak minimisation condition:
$\frac{m_{Z}^{2}}{2}=\frac{(m_{H_{d}}^{2}+\Sigma_{d}^{d})-(m_{H_{u}}^{2}+\Sigma_{u}^{u})\tan^{2}\beta}{(\tan^{2}\beta-1)}-\mu^{2}\simeq-m_{H_{u}}^{2}-\mu^{2}.$ (7)
where the $\Sigma_{u,d}^{u,d}$ contain a large assortment of SUSY loop corrections. Barbieri and Giudice (1988) introduced the naturalness measure
$\Delta_{\rm BG}\equiv max_{i}\left|\frac{\partial\log m_{Z}^{2}}{\partial\log p_{i}}\right|\,,$ (8)
where the $p_{i}$ are fundamental parameters of the theory, and used this to set bounds on SUSY masses by imposing a maximally allowed value of $\Delta_{\rm BG}^{\rm max}=10$. Following this idea, Baer et al. (2016) formulated a naturalness principle stating that “An observable ${\cal O}$ is natural if all independent contributions to ${\cal O}$ are less than or of order ${\cal O}$”. This is dubbed “practical naturalness”. Applied to eq. (7), practical naturalness requires each term on the right-hand-side (RHS) of the equation to be $\lesssim m_{Z}^{2}/2$. This is used to define an electroweak fine-tuning measure $\Delta_{\rm EW}$ as
[…]
Note that $\Delta_{\rm BG}$ agrees with $\Delta_{\rm EW}$ for both, weak-scale and high-scale inputs, when evaluated in terms of genuinely independent parameters. Moreover, the naturalness problem related to $m_{h}^{2}$ evoked in the previous sections becomes much less severe in terms of practical naturalness.
Baer et al. (2016) conclude that gluinos, top squarks and other SUSY particles can live at the multi-TeV scale with little cost to naturalness (because they have loop-suppressed contributions to the weak scale), but higgsinos, whose masses are essentially given by $\mu$, must lie close to the weak scale, with $\mu\lesssim 200-400$ GeV. As we will see below, such light higgsinos have a tendency to escape detection at the LHC. Even at the High-Luminosity LHC, much but not all of the $\mu\lesssim 200-400$ GeV parameter space can be covered, see Baer et al. (2020).
Figure 4 shows the mass ranges compatible with $\Delta_{\rm EW}<30$ in the Non-Universal Higgs Mass Model (NUHM2). Also shown are the upper bounds withing the phenomenological MSSM (pMSSM), which is the topic of the next section.
Figure 4: Range of sparticle and heavy Higgs masses allowed by naturalness within the NUHM2 model with $\Delta_{\rm EW}<30$. Also imposed are the LEP2 limit on charginos, LHC Run 1 bounds on gluino and squark masses, and $m_{h}=125\pm 2$ GeV. The black bars show upper bounds from the pMSSM model with 19 weak scale parameters; from Baer et al. (2016).
4 Phenomenological MSSM
The pMSSM is a 19-dimensional realisation of the R-parity conserving MSSM with parameters defined at the SUSY scale $M_{S}=\sqrt{m_{\tilde{t}_{1}}m_{\tilde{t}_{2}}}$. It was introduced in Djouadi et al. (1998) to provide a more general framework for studying MSSM phenomenology than the predominant minimal or constrained SUSY-breaking models, without having to deal with the $>100$ free parameters of the general MSSM, which would be unwieldy for most phenomenological purposes.
Figure 5: Fraction of pMSSM points that are not excluded at 95% confidence level by ATLAS SUSY searches from Run 2 (given in Table 4 of Arbey et al. (2022)), and that remain beyond the projected sensitivity of Run 3, shown as a function of the gluino mass (upper panel) and the lighter stop mass (lower panel); from Arbey et al. (2022).
The pMSSM employs only a few plausible assumptions motivated by experiment. Specifically, all soft SUSY-breaking parameters are taken to be real to avoid introducing new CP-violating effects. The sfermion mass matrices and trilinear coupling matrices are assumed to be diagonal, preventing FCNCs at tree level. Additionally, the first and second generations of sfermions are treated as degenerate at low energy, and trilinear couplings are assumed to be the same for all three generations. This leaves the pMSSM with 19 free parameters.
This setup allows for generic mass and mixing patterns, leading to a much larger variety of cross sections, decay channels and branching ratios, and thus to a much broader set of predictions for experimental signatures than those found in any of the constrained scenarios, such as the CMSSM, mSUGRA, NUHM models, etc., which impose quite restrictive relations between the soft SUSY-breaking parameters (see AbdusSalam et al. (2010); Berger et al. (2009); Conley et al. (2011); Sekmen et al. (2012); Arbey et al. (2012b) for pioneering studies). Likewise, it also leads to a much broader set of predictions than realised in simplified models.
Consequently, a significant portion of the pMSSM parameter space escapes the bounds from ATLAS and CMS searches, which were set within constrained or simplified models. This was confirmed by the ATLAS and CMS reinterpretations of their Run 1 searches, see Aad et al. (2015); Khachatryan et al. (2016). Indeed, mass limits derived within simplified models are often overly optimistic and do not hold in the more general case of the pMSSM (or other realistic SUSY models).
As an example, Fig. 5 shows the fraction of pMSSM points (after flavour, dark matter, and low-energy constraints obtained using the SuperIso program Mahmoudi (2008, 2009)) not excluded by ATLAS Run 2 searches as a function of the gluino and stop masses. In the context of simplified models, the considered ATLAS analyses (given in Table 4 of Arbey et al. (2022)) exclude gluino masses up to $2.2-2.35$ TeV depending on the LSP mass, and stop masses in the range $450-950$ GeV for LSP masses below 160 GeV (in the case where $m_{\tilde{t}_{1}}\sim m_{t}+m_{\tilde{\chi}^{0}_{1}}$, stops in the range of $235-590$ GeV are excluded). This is to be compared with the full lines in Fig. 5, which indicate that gluinos with masses as low as 1 TeV and stops as light as 400 GeV remain viable within the pMSSM framework. [^4] It should be noted that the figure is not intended to provide a statistical interpretation; rather, it serves to illustrate that viable spectra with relatively light superpartner masses are indeed allowed within the pMSSM framework. Also shown in Fig. 5 are the projected sensitivities of Run 3 with 400/fb (dashed lines). As one can see, the fraction of surviving points is reduced, but the overall conclusion remains unchanged.
Turning to electroweak-inos (i.e. charginos and neutralinos, mixtures of electroweak gauginos and higgsinos), in Aad et al. (2024) the ATLAS collaboration provided an interpretation of their Run 2 searches for electroweak production of supersymmetric particles within the pMSSM. To this end, they scanned bino, wino and higgsino mass parameters $M_{1}$, $M_{2}$ and $\mu$ from $-2$ TeV to $2$ TeV; the gluino mass parameter $M_{3}=1-5$ TeV; $m_{A}=0-5$ TeV and $\tan\beta=1-60$. The masses of 1st/2nd generation sfermions were fixed to 10 TeV and the 3rd generation parameters scanned over to obtain scenarios with a SM-like Higgs in the right mass range. Considering the results from eight separate searches for electroweak-inos (see Table 5 in Aad et al. (2024)), each using 140/fb of proton-proton data from Run 2, it was found that many pMSSM scenarios with chargino and neutralino masses well below the simplified-model limits escape exclusion, as examplified in Fig. 6.
Figure 6: Fraction of electroweak-ino scenarios excluded by ATLAS Run 2 results. Considered are 7956 pMSSM points that have a neutralino LSP, satisfy chargino mass constraint from LEP and have a light Higgs mass in the range $m_{h}=120-130$ GeV. The white line shows the limit in the simplified wino/bino model; from Aad et al. (2024).
Apparently, in the pMSSM, the total SUSY signal tends to spread out over several experimental channels. It is therefore interesting to ask how the sensitivity improves when statistically combining the likelihoods from different, independent analyses. Likewise, it is interesting to combine ATLAS and CMS results, to increase statistics and mitigate fluctuations in the data. This is possible using the SModelS framework as demonstrated recently in Altakach et al. (2023, 2024b). Since Aad et al. (2024) made all their scan points publicly available, we can use SModelS to reevaluate the constraints in a global likelihood analysis. Concretely, we are combining up to 33 analyses [^5] The best (i.e. most sensitive) combination is determined for each parameter point separately. Only analyses that can to good approximation be considered as uncorrelated enter the combination. from both Run 1 and Run 2 (21 from ATLAS and 12 from CMS) implemented in the SModelS v3.0.0 database [Altakach et al. (2024a)] that are sensitive to the electroweak-ino pMSSM scan points.
The result is shown in Fig. 7. The plot shows (in colour) the allowed points, i.e. the points not excluded at 95% confidence level by the global likelihood, in the plane of lightest neutralino ($\tilde{\chi}^{0}_{1}$) versus lightest chargino ($\tilde{\chi}^{\pm}_{1}$) mass. The colour code indicates the combined observed $r$-value, $r^{\rm comb}_{\rm obs}$, with $r$ defined as the ratio of the predicted fiducial cross section of the signal over the corresponding upper limit. [^6] Points allowed by the combination but excluded in Aad et al. (2024) by the disappearing tracks analysis with 140/fb, for which only the 36/fb results are available in SModelS, have been removed. The grey points have $r^{\rm comb}_{\rm obs}\geq 1$ and are thus excluded by the combination.
As one can see, only a small region of the pMSSM parameter space is definitely excluded, while many scenarios with light charginos and neutralinos remain valid. The colour code in Fig. 7 gives an idea about future prospects: very roughly, the red/orange/yellow-ish points should be within reach with more data at Run 3, while the blue(-ish) points will be hard to test (pending the excesses discussed in the next section).
Figure 7: Scatter plot of pMSSM scan points from Aad et al. (2024) that escape exclusion in a global likelihood analysis with SModelS (in colour). Grey points are excluded by the combination of analyses. See text for details.
When requiring $|\mu|<400$ GeV, as preferred by the naturalness argument elaborated on in the previous section, only the region with $m_{\tilde{\chi}^{\pm}_{1}}<400$ GeV remains from Fig. 7, with the density of points considerably thinned out, see Fig. 8.777The low density of points with $|\mu|<400$ GeV is simply a volume effect from the sampling of a vast parameter space. The overall picture, however, remains unchanged, with many well-motivated light SUSY scenarios escaping current limits. Specifically one finds along the diagonal many higgsino LSP points with very low $r^{\rm comb}_{\rm obs}$, which may be challenging to detect also in future LHC runs; these typically have heavy winos ($|M_{2}|>1$ TeV). The higgsino LSP points with a promising $r^{\rm comb}_{\rm obs}$ are typically accompanied by winos below about 1 TeV. There are also points with wino-like LSPs, but with a large enough higgsino or bino admixture to make the $\tilde{\chi}^{\pm}_{1}$ not too long-lived (or even decay promptly). As pointed out in Carpenter et al. (2024), part of this parameter space could be tested by hadronic mono-boson searches, in particular if the MET cuts are optimised. Away from the diagonal, that is for sizeable $\tilde{\chi}^{\pm}_{1}-\tilde{\chi}^{0}_{1}$ mass differences, the spectrum consists of light higgsinos with a bino LSP; again the $r$-values are higher when winos are not too heavy, so they also contribute to the signal. Details will be discussed in Constantin et al. (2025).
Figure 8: Similar to Fig. 7 but requiring $|\mu|<400$ GeV.
5 Hints for light electroweak-inos?
Last but not least, we note that there are tantalising, small but seemingly consistent, excesses in ATLAS and CMS supersymmetry searches involving two or more soft leptons, see Aad et al. (2020, 2021a); Tumasyan et al. (2022); Hayrapetyan et al. (2024). The excesses mostly appear in the dilepton ($2\ell$+MET) channel, with a dilepton invariant mass in the range of $10-20$ GeV, and can be interpreted as stemming from light charginos/neutralinos, $pp\to\tilde{\chi}^{\pm}_{1}\tilde{\chi}^{0}_{2}\to W^{*}\tilde{\chi}^{0}_{1}Z^{*}\tilde{\chi}^{0}_{1}$, with a compressed mass spectrum, roughly around $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 10-30$ GeV. Small excesses are also seen in monojet searches [Aad et al. (2021b); Tumasyan et al. (2021)] and, as pointed out in Agin et al. (2024a), may be compatible with the soft lepton ones.
The simplified models used by ATLAS and CMS for their interpretations, a wino/bino model with $m_{\tilde{\chi}^{0}_{2}}=m_{\tilde{\chi}^{\pm}_{1}}$ and a higgsino LSP model with $m_{\tilde{\chi}^{\pm}_{1}}=0.5(m_{\tilde{\chi}^{0}_{2}}+m_{\tilde{\chi}^{0}_{1}})$, are, however, never exactly realised in the MSSM. In the wino/bino case, for the small mass splittings relevant here, there will always be some wino-bino mixing. Moreover, as pointed out by Baum et al. (2023); Chakraborti et al. (2024), $\tilde{\chi}^{0}_{2}\to\tilde{\chi}^{0}_{1}\gamma$ decays may be relevant and weaken the leptonic signal. In the higgsino LSP case, on the other hand, a non-negligible gaugino admixture is needed to make the mass splitting large enough, as pure higgsinos have loop-suppressed mass splittings below about 1 GeV.
Realistic MSSM scenarios for explaining the observed excesses, in part including also dark matter (DM) constraints, were considered in Agin et al. (2024a); Chakraborti et al. (2024); Agin et al. (2024b); Martin (2024). Chakraborti et al. (2024) concluded that wino/bino DM in the MSSM performs better than the higgsino scenario (for $\mu>0$) and for $m_{\tilde{\chi}^{0}_{2}}\approx m_{\tilde{\chi}^{\pm}_{1}}\leq 250$ GeV and $\Delta m=m_{\tilde{\chi}^{0}_{2}}-m_{\tilde{\chi}^{0}_{1}}\approx 25$ GeV, has roughly the right cross section to produce the observed excesses. Martin (2024) considered either sign of $\mu$ and concluded that $\mu<0$ and low $\tan\beta$ (with decoupled heavy Higgs bosons) allows one to evade DM direct detection limits, rendering light higgsinos a viable option. Agin et al. (2024b) discussed the complementarity of the soft lepton excesses with other excesses in LHC monojet searches for the MSSM and next-to-MSSM, as well as for two non-SUSY scenarios (a vector-like lepton model and a type-II see-saw model); they also identified the wino/bino scenario in the MSSM as “the best option currently on the table”, while the non-SUSY interpretations fit the excesses less well.
In any case, more data is needed to assess whether the current small excesses are indeed hints of new physics or mere fluctuations. Run 3 should clarify the situation.
6 Summary and conclusions
Has the LHC ruled out supersymmetry? The answer is no!
As we have shown, interesting and well-motivated scenarios with light superpartners remain viable even in the simplest incarnation of SUSY, the MSSM. Discovery prospects at the LHC remain good, and there are even possible hints for light electroweak-inos in the current data. However, since High-Scale SUSY is also a viable possibility, no currently feasible collider can definitely exclude SUSY as a principle of nature.
On the other hand, as discussed in, e.g., Dumont et al. (2014); Arbey et al. (2016, 2022), precise measurements of the Higgs boson offer a valuable and complementary probe of supersymmetry. The presence of additional SUSY particles can modify the couplings of the SM-like Higgs, thereby altering its production cross sections and decay branching ratios. As a result, precision Higgs measurements can impose meaningful constraints on the SUSY parameter space, even in scenarios where direct searches are not effective.
All in all, the hunt for SUSY remains exciting.
Acknowledgements
We thank the editor for the invitation to write this review.